OpenStudy (anonymous):

Solve for x: 1000 = 500e^2x a. x=ln(2) / 2 b. x=ln(1000) / 2*ln(500) c. x=1/e d. x=ln(2)-2

OpenStudy (anonymous):

First thing to do is get rid of the \(500\). Then you can use the natural logarithm to get rid of the \(e\).

OpenStudy (anonymous):

Can you help me?

OpenStudy (anonymous):

1000 = e^2x what do I do now?

OpenStudy (marigirl):

1000 = 500e^2x divide both sides by 500 2= e^2x take the log of both sides ln(2)=2x (*Note ln and e cancel out leaving you with 2x) . Divide both sides by 2 x= (ln2)/2

OpenStudy (anonymous):

oh I understand. Thank you